OpenStudy (anonymous):

Celine has a bottle that contains 20% of milk and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle. Define the variable in the equation. Part B: How many liters of milk are present in the bottle? Show your work. I got Part A already: b=.20b+1, b represents what the bottle contains entirely. Part B I need help with though. I don't know how to solve b=.20b+1

OpenStudy (anonymous):

The problem states that only one variable can be used. Let L be the total number of liters of water and milk combined. 20% of the total is Milk. Therefore, Milk = L x 20/100 = 0.2 L The rest of it is Water. So Water = L - 0.2L = 0.8L The problem states the bottle has 1 liter of water. So 0.8L = 1 Answer to a) 0.8L = 1 is the equation where L is the total number of liters of water and milk combined. Milk = 0.2L and Water = 0.8L b) Solve for L in 0.8L = 1 L = 1/0.8 = 1.25 Milk = 0.2L = (0.2)(1.25) = 0.25 liters So this what I think, hope i helped :)

OpenStudy (anonymous):

v milk v water v bottle 0.2 * L + 0.8*L = 1 for milk: 1 - 0.8*L = milk for water: 1 -0.2*L = water

OpenStudy (anonymous):

@mathessentials, could you show me how to get it?